OpenStudy (livya15):

This is a Calculus question: I need help understanding algebraic functions. Function given below. Things in need to find: domain; points of discontinuity and type. Please Help! Test tomorrow!!!

OpenStudy (livya15):

(Square root) of -/x^2 -9 Divided by. X+3

OpenStudy (anonymous):

i am going to guess it is \[f(x)=\frac{\sqrt{x^2-9}}{x+3}\] but i could be wrong

OpenStudy (livya15):

Correct thank you...that thing wouldn't come up for me :)

OpenStudy (livya15):

@satellite73

OpenStudy (livya15):

@Hero

OpenStudy (anonymous):

ok domain first

OpenStudy (anonymous):

it should be pretty clear right away before we worry about the numerator that \(x\neq -3\) right?

OpenStudy (livya15):

Alright

OpenStudy (anonymous):

and since you cannot take the square root of a negative number, we also have to solve \[x^2-9\geq 0\] as well do you know how to solve that for \(x\) ?

OpenStudy (livya15):

(X-3)(x+3)>(or equal to) 0

OpenStudy (anonymous):

xactly you know the zeros are at -3 and 3, and since it is a parabola that opens up (i mean \(y=x^2-9\) is the parabola) then it must be positive outsides the zeros and negative between them you want 'positive" so it will be \[(-\infty,-3]\cup [3,\infty)\]but now we have to be careful

OpenStudy (anonymous):

since we have excluded \(-3\) from the domain, the real answer is \[(-\infty, -3)\cup [3,\infty)\]

OpenStudy (livya15):

Could you write that answer in another way ?

OpenStudy (anonymous):

sure \(x<-3\) or \(x\geq 3\)

OpenStudy (livya15):

Ok :)

OpenStudy (livya15):

So that is the domain, so how do you find the points of discontinuity?

OpenStudy (anonymous):

it is discontinuous at the end points of the domain

OpenStudy (anonymous):

at 3 it is discontinuous because the function does not exist to the left of 3

OpenStudy (anonymous):

at -3 it is discontinuous for the same reason, but also because as x goes to -3 from the left, the whole thing goes to \(-\infty\)

OpenStudy (livya15):

3 would be a nonremovable point of discon. And -3 would be a removable pt. of discon.?

OpenStudy (livya15):

@satellite73

OpenStudy (anonymous):

neither are removable

OpenStudy (anonymous):

removable basically means you have a hole in the graph you can mostly remove it by factoring and cancelling to get an equivalent expression that is not true in this case

OpenStudy (livya15):

So the lim f(x)= -infinity X->3- Lim f(x)= +infinity X->3+

OpenStudy (anonymous):

first one is true second one is not

OpenStudy (anonymous):

\[\lim_{x\to 3^+}f(x)=0\]

OpenStudy (anonymous):

then numerator goes to zero, the denominator goes to 6, and \(\frac{0}{6}=0\)

OpenStudy (livya15):

Oh ok! Thank you!,

OpenStudy (anonymous):

yw